OpenStudy (anonymous):

The number who attended this year's conference was 2 3/8 times the number who attended the conference last year. if 4180 people attended this years conference, how many attended last year?

OpenStudy (anonymous):

2+3/8=2.375 x*2.375=4180 x=4180/2.375 x= 1760

OpenStudy (anonymous):

You write down the given by translating it with an equation: 2(3/8) of (x) people attended last year = 4180 people from this year (was means = ) so: solve for x and you'll get: x = 5573 people so approximately 5573 people attended Please correct me if I'm wrong ^_^